# “Feeling that I was Collaborating with Them:” A 20-year Systematic Literature Review of Social Virtual Reality Leveraging Collaboration

### v2

@@ 2.3. Collaboration Work in Social VR
Collaboration in Social VR refers to two or more individuals working together within a Social VR application to achieve a common goal (Freeman et al., 2022). The distinct affordances provided by Social VR, such as avatar customization, nuanced non-verbal behaviors, and spatial interactions, can facilitate a strong sense of being together and foster relationship building among distributed collaborators (Freeman and Acena, 2021). It also enables remote teams to engage in a wide range of collaborative activities that mirror in-person teamwork, from creative projects and problem-solving to mundane tasks and social experiences (Freeman et al., 2022). Furthermore, Social VR allows the interplay of productivity and sociality that often characterizes co-located collaboration (Maloney and Freeman, 2020). In one study, Maloney et al. (Maloney and Freeman, 2020) conducted interviews to explore what makes activities on Social VR platforms meaningful to their users. They found highlighted immersive interactions through full-body tracking and facilitating more natural and expressive communication as the main advantages of Social VR, overcoming the limitations of 2D video conferencing. In another study, Li et al. (Li et al., 2021) argue that Social VR affords more natural social interactions than video conferencing, such as the ability for users to organically break off into small groups and interact with virtual objects in the shared environment. Unlike video conferencing, which restricts users to ’talking head experiences’—static, front-facing video feeds—Social VR allows for more physical activities and spontaneous collaborations that naturally arise from users’ interactions (Li et al., 2021).
@@
Given the growing body of research on Social VR, we aim to systematically map how previous research has studied these technologies in the context of collaborations and teamwork. As the CSCW community has delved into several questions and reflections on Social VR, synthesizing key findings and implications from these studies can bring new opportunities for understanding the specific benefits of Social VR for team dynamics, designing more effective collaboration applications, and understanding the barriers to adoption. As technologies continue to shape how people work together (Harris et al., 2019), the CSCW community is well positioned to examine how Social VR’s unique capabilities and affordances enable remote collaborations and guide the development of future technological ecosystems (Wallace et al., 2017; Gonzales et al., 2015).
@@ 3. Methodology
We followed a scoping review methodology (Arksey and O’Malley, 2005) to synthesize existing knowledge and reveal trends in using Social VR for collaboration purposes. This methodology allows researchers to map the literature and identify gaps in a specific area of research (Arksey and O’Malley, 2005; Levac et al., 2010). Scoping reviews have several advantages: they are transparent, comprehensive, less prone to bias, and make it easier to reproduce the detailed information reported about each step of the review and how it is conducted (Pham et al., 2014). They have gained prominence in the HCI and CSCW communities as they allow for the synthesis and comprehension of the development of different research directions (Wallace et al., 2017; Cosio et al., 2023; Bergram et al., 2022; Rogers et al., 2022; Wei et al., 2022). We used the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) guidelines to report the methods and results (Page et al., 2021). PRISMA covers four stages in the review process: Identification, Screening, Eligibility, and Inclusion. We collected papers and tabulated them in a shared Google Spreadsheet, capturing metadata such as publication year, abstract, and keywords. Once the metadata was recorded, the authors of the articles were hidden to avoid potential bias during the coding phase.
@@ 3.1. Search Strategy and Data Sources
To identify relevant articles for our systematic literature review, we developed a comprehensive search query that could capture the key concepts of our study while balancing sensitivity and specificity. After several iterations, we built our search query by breaking down our research topic into its core components: the technology (i.e., social VR), the activity (i.e., collaboration), and the groups (i.e., teams). For the technology concept, we included several synonyms and related terms to maximize coverage, namely “Social Virtual Reality,” “Social VR,” “Multi-user Virtual Reality,” and “Collaborative Virtual Environments.” We intentionally did not include the term “Virtual Reality” alone, as our focus was on Social VR environments specifically designed to enable interaction and collaboration rather than VR more broadly. For the collaboration concept, we chose a set of keywords focused on people working together: “Collaborative,” “Collaboration,” and “Cooperation.” We chose to add the term “cooperation” alongside “collaboration,” as the two are closely related and often used interchangeably in the literature (Stout and Keast, 2021). Lastly, for the groups concept, we included the words referring to teams: “Team,” “Teams,” “Teamwork,” “Group,” and “Groups.” These words defined the importance of studies analyzing multiple individuals or users within a unit. We combined the synonyms within each concept using the Boolean OR operator and then linked the three main concepts with the AND operator. As a result, our final query was: (“Social Virtual Reality” OR “Social VR” OR “Multi-user Virtual Reality” OR “Collaborative Virtual Environments”) AND (Team OR Teams OR Teamwork OR Group OR Groups) AND (Collaborative OR Collaboration OR Cooperation).
@@
We conducted the first step, “Identification,” using the following three databases: IEEE Xplore Digital Library, ACM Digital Library, and Web of Science. We chose these datasets since they cover the leading conferences and journals on human-computer interaction and team science. We searched for articles in these databases based on their titles, abstracts, and full text. We conducted the final search on February 19, 2025. We included articles published between 2004 and 2024, covering 20 years of research. We limited the included papers to the last 20 years to ensure that our corpus captured the most recent developments, which led to the significant boom in HMD technologies starting around the 2010s, including HTC Vive and Oculus Rift. We also checked earlier studies, which began as early as 1993 in our search scope, and primarily focused on collaboration in virtual environments using screen-based technologies rather than HMDs. For the few early papers involving VR, their contributions were either revisited in later research or shaped by the technological constraints of the time, which limited their applicability to the current landscape of Social VR research. We obtained 892 papers from the IEEE Xplore Digital Library, 972 papers from the ACM Digital Library, and 182 papers from the Web of Science database, resulting in a total of 2,046 papers. Since the ACM Digital Library did not allow users to export the results from their website, we manually copied and pasted the ACM Digital Library’s results to a Google Spreadsheet. We exported the results of all three databases from their websites to a CSV file. We then compiled all the results into a single spreadsheet for analysis. In cases where we found two versions of an article (e.g., a conference paper later extended into a journal article), we only included the most recent version. Based on the titles, we removed 11 duplicate articles. This left us with 2,035 articles for screening. Fig. 2 depicts the trend of paper counts for each year in our search corpus from 2004 to 2024.
@@ 3.3. Article Selection
[…]
PRISMA Flow Diagram of this study. It presents the details of the article selection process.
@@ 3.4. Data Extraction and Synthesis
[…]
The number of included articles in three databases between 2004-2024.
@@ 3.5. Thematic Analysis
We conducted a thematic analysis to identify cohesive themes that captured patterns across the final corpus (Braun and Clarke, 2006). The goal of this analysis was to inductively derive themes grounded in the data while minimizing confirmation bias. The coder re-reviewed the articles, including their extracted data and the notes summarizing the main contributions, to generate initial codes. These codes were developed inductively, without reference to pre-existing coding frameworks or theoretical models (Nowell et al., 2017; Braun and Clarke, 2006). The codes were not mutually exclusive since the articles could address multiple codes. The coder iterated the initial codes through multiple rounds of analysis until no new ideas emerged from them. Following this, the refined codes were grouped into themes, which were further revised and renamed over several rounds to improve their clarity and alignment with the corpus data. To enhance reliability and interpretive consensus, all the authors reviewed and discussed the themes and their boundaries in multiple meetings. Ultimately, the final themes were organized into three overarching categories and nine subcategories that captured the breadth and nuance of how Social VR has been studied in the context of collaboration.
@@ 4. Results
After applying the methodology described above, 62 of the 2,035 papers were included in our final corpus. Fig.  3 shows the filtering process of the papers in IEEE Xplore Library, ACM Digital Library, and Web of Science databases. The most common reason for excluding articles during the eligibility stage was that they were not related to collaboration in VR—meaning the study either focused on collaboration without using VR (e.g., in 2D platforms), used VR without addressing collaboration, or addressed neither. This exclusion category is labeled in Fig.  3 as ”not related to VR, collaboration, or both” (n = 356). Following this process, Fig. 4 depicts the trend in the number of papers for each year among the 62 included articles from 2004 to 2024.
@@ 4.1. Description of the Included Articles
Table 1 provides an overview of research and experimental characteristics of the 62 included articles. Table 1a summarizes the research types, evaluation methodologies, and analysis methods. The majority were laboratory experiments (69.35%), followed by non-empirical papers (12.90%), survey papers (8.07%), and qualitative descriptive studies (4.84%). Three articles (4.84%) fell under other categories, including one case study, one conceptual framework, and one system architecture description.
Regarding evaluation methodology, 33.87% of the studies employed mixed methods, 32.26% used quantitative methods, and 17.74% used qualitative approaches. In 16.13% of the articles, an evaluation methodology was not employed or specified. Among the qualitative methods articles, thematic analysis was the most common method (19.36%), followed by free-text responses (4.84%) and participant experience summaries (3.23%). A subset of studies employed methods such as open coding, observational techniques, or participant experience summaries. On the quantitative side, ANOVA was the most frequently used analysis (19.28%), followed by descriptive statistics (15.67%), t-tests and Wilcoxon tests (12.05% each), and post-hoc or Mann-Whitney U tests (combined 13.24%). Since many studies applied more than one technique, the total number of quantitative methods exceeds the number of papers.
@@
In terms of types of tasks, problem-solving was most common task employed in these articles (23.26%), followed by navigation/ wayfinding (16.28%) and design/construction (16.28%). Other tasks included negotiation/planning (11.62%), meetings/discussions and data analysis (9.30% each), scientific tasks (6.98%), and one workplace simulation (2.33%). Two studies (4.65%) included task types that did not fall under the primary categories outlined in the table.
Lastly, most articles with laboratory studies focused on small teams, including dyads (two-person groups, 79.08%) and triads (16.28%). Only two studies involved more participants per group, one study with four participants, and another with seven participants.
@@ 4.2. Thematic Analysis
The first theme includes articles addressing the foundational aspects of individual perception and experience within Social VR environments. This theme includes two sub-themes: “Am I Really There?,” which highlights how Social VR can simulate users’ physical presence, enhancing their interaction and engagement with the collaborative task. The second sub-theme, “How Am I Perceived Here?,” explores avatar design and its impact on collaborative dynamics and engagement, emphasizing the importance of avatar customization to shape user interactions and self-expression within virtual spaces.
@@
The final theme examines how collaboration is supported by the unique affordances of Social VR. These affordances refer to the technical elements and interactions designed to enable realistic and functional 3D virtual environments. This theme includes three sub-themes: “Creating Realistic and Functional Virtual Environments,” which addresses how the design of virtual spaces that mimic real-world settings can improve user experience; “Facilitating Interaction through Virtual Tools and Objects,” which emphasizes the importance of object manipulation and tool use in collaborative tasks; and “Enhancing Immersion through Spatial Audio and Movement,” which underscores how spatial audio and coordinated movement contribute to a more immersive and effective collaborative experience.
A summary of these themes, sub-themes, and their respective articles can be found in Table 2.
@@
“Am I Really There?” Sense of Presence as a Foundation for Immersive Collaboration. 4 (Oh et al., 2018), (Olaosebikan et al., 2022), (Khojasteh and Won, 2021), (Mei et al., 2021)
“How Am I Perceived Here?” Avatar Design Shaping Collaborative Dynamics and Engagement. 7 (Osborne et al., 2023), (Torro and Pirkkalainen, 2023), (Yang et al., 2024), (Yassien et al., 2021), (Freeman et al., 2020), (Kimmel et al., 2024), (Kocur et al., 2020)
@@
“Are We Really Together?” Social Presence Enhancing Team Collaboration. 7 (Olaosebikan et al., 2022), (Dey et al., 2024), (Darbar et al., 2024), (Bovo et al., 2022), (Ide et al., 2020), (Adkins et al., 2024), (Mei et al., 2021)
@@
“Are We Fully Engaged?” The Importance of Social Engagement in VR Collaboration. 3 (Xu et al., 2023), (Le Tarnec et al., 2024), (Moharana et al., 2023)
@@
Creating Realistic and Functional Virtual Environments. 6 (Mei et al., 2021), (Osborne et al., 2023), (Men and Bryan-Kinns, 2019), (Freiwald et al., 2020), (Weissker and Froehlich, 2021), (Langa et al., 2022)
Facilitating Interaction through Virtual Tools and Objects. 4 (García et al., 2008), (Dey et al., 2024), (Osborne et al., 2023), (Roberts et al., 2005)
Enhancing Immersion through Spatial Audio and Movement. 6 (Weissker et al., 2019), (Immohr et al., 2024), (Ríos et al., 2018), (Immohr et al., 2023), (Weissker, 2020), (Naef et al., 2004)
Table 2. Themes and sub-themes identified through thematic analysis of Social VR research on collaboration. Each row presents a sub-theme, the number of articles (N) associated with it, and its references.
[…]
Themes and sub-themes identified through thematic analysis of Social VR research on collaboration. Each row presents a subtheme, the number of articles (N) associated with it, and representative references. The themes are grouped into three overarching categories reflecting different levels of collaborative experience in Social VR: individual perception, team dynamics, and technological affordances.

### v3

@@ 2.3. Collaboration Work in Social VR
Collaboration in Social VR refers to two or more individuals working together in a Social VR application toward a common goal (Freeman et al., 2022). The distinct affordances provided by Social VR, such as avatar customization, nuanced non-verbal behaviors, and spatial interactions, can facilitate a strong sense of being together and foster relationship building among distributed collaborators (Freeman and Acena, 2021). It also enables remote teams to engage in various collaborative tasks that mirror in-person teamwork, from creative projects and problem-solving to mundane tasks and social experiences (Freeman et al., 2022). Furthermore, Social VR allows the interplay of productivity and sociality that often characterizes co-located collaboration (Maloney and Freeman, 2020). In one study, Maloney et al. (Maloney and Freeman, 2020) conducted interviews to explore what makes activities on Social VR platforms meaningful to their users. They found interactions through full-body tracking and facilitating more natural and expressive communication as the main advantages of Social VR, overcoming the limitations of 2D video conferencing. In another study, Li et al. (Li et al., 2021) argue that Social VR affords more natural social interactions than video conferencing, such as the ability for users to organically break off into small groups and interact with virtual objects in the shared environment. Unlike video conferencing, which restricts users to ‘talking head experiences’—static, front-facing video feeds—Social VR allows for more physical activities and spontaneous collaborations that naturally arise from users’ interactions (Li et al., 2021).
@@
Given the growing research on Social VR, we aim to map how previous research has studied these technologies in the context of collaborations and teamwork. Synthesizing key findings and implications from these studies can reveal the specific benefits of Social VR for team dynamics, inform the design of more effective and safe collaboration applications, and identify barriers to adoption. As technologies continue to reshape how people work together (Harris et al., 2019), the CSCW community is well positioned to examine how Social VR’s unique capabilities and affordances enable remote collaborations and guide future system development (Wallace et al., 2017; Gonzales et al., 2015).
@@ 3. Methodology
We followed a scoping review methodology (Arksey and O’Malley, 2005) to synthesize existing knowledge and identify trends in using Social VR for collaborative purposes. This methodology allows researchers to map the literature and identify gaps in a specific area of research (Arksey and O’Malley, 2005; Levac et al., 2010). Scoping reviews have several advantages: they are transparent, comprehensive, less prone to bias, and make it easier to reproduce the detailed information reported about each step of the review and how it is conducted (Pham et al., 2014). They have gained prominence in the HCI and CSCW communities as they enable the synthesis and comprehension of the development of different research directions (Wallace et al., 2017; Cosio et al., 2023; Bergram et al., 2022; Rogers et al., 2022; Wei et al., 2022).
We used the Preferred Reporting Items for Scoping Reviews and Meta-Analyses (PRISMA) guidelines to report the methods and results (Page et al., 2021). PRISMA covers four stages in the review process: Identification, Screening, Eligibility, and Inclusion. We collected papers and tabulated them in a shared Google Spreadsheet, capturing metadata such as publication year, abstract, and keywords. Once the metadata was recorded, authors’ names were hidden to avoid bias during the coding phase.
Consistent with scoping-review aims, we did not conduct a formal critical appraisal of individual studies; our goal was to map the evidence and characterize study features.
@@ 3.1. Search Strategy and Data Sources
To identify relevant articles for our scoping review, we developed a comprehensive search query that could capture the key concepts of our study while balancing sensitivity and specificity. After several iterations, we built our search query by breaking down our research topic into its core components: the technology (i.e., social VR), the activity (i.e., collaboration), and the groups (i.e., teams). For the technology concept, we included several synonyms and related terms to maximize coverage, namely “Social Virtual Reality,” “Social VR,” “Multi-user Virtual Reality,” and “Collaborative Virtual Environments.” We intentionally did not include the term “Virtual Reality” alone to focus on studies that specifically addressed Social VR. For the collaboration concept, we chose a set of keywords focused on people working together: “Collaborative,” “Collaboration,” and “Cooperation.” We chose to add the term “cooperation” alongside “collaboration,” as the two are closely related and often used interchangeably in the literature (Stout and Keast, 2021). Lastly, for the groups concept, we included the words referring to teams: “Team,” “Teams,” “Teamwork,” “Group,” and “Groups.” These words defined the importance of studies analyzing multiple individuals or users within a unit. We combined the synonyms within each concept using the Boolean OR operator and then linked the three main concepts with the AND operator. As a result, our final query was: (“Social Virtual Reality” OR “Social VR” OR “Multi-user Virtual Reality” OR “Collaborative Virtual Environments”) AND (Team OR Teams OR Teamwork OR Group OR Groups) AND (Collaborative OR Collaboration OR Cooperation).
@@
We conducted the first step, “Identification,” using the following three databases: IEEE Xplore Digital Library, ACM Digital Library, and Web of Science. We chose these datasets since they cover the leading conferences and journals on human-computer interaction and team science. We searched for articles in these databases using their titles, abstracts, and full text. We conducted the final search on February 19, 2025. We included articles published between 2004 and 2024, covering 20 years of research. We limited the included papers to the last 20 years to ensure that our corpus captured the most recent developments, which led to the significant boom in HMD technologies starting around the 2010s, including HTC Vive and Oculus Rift. We also checked earlier studies, which began as early as 1993 in our search scope, and primarily focused on collaborative virtual environments (CVEs) (Benford et al., 2001; Churchill and Snowdon, 1998) using screen-based technologies rather than HMDs. For the few early papers involving VR, their contributions were either revisited in later research or shaped by the technological constraints of the time, which limited their applicability to the current landscape of Social VR research.
We obtained 892 papers from the IEEE Xplore Digital Library, 972 papers from the ACM Digital Library, and 182 papers from the Web of Science database, totaling 2,046 papers. Since the ACM Digital Library does not allow users to export results from its website, we manually copied and pasted the results into a Google Spreadsheet. For the other two datasets, we exported the results from their websites to a CSV file. We then compiled all the results into a single spreadsheet for analysis. In cases where we found two versions of an article (e.g., a conference paper later extended into a journal article), we included the most recent version. Based on the titles, we removed 11 duplicate articles. This left us with 2,035 articles for screening. Fig. 2 depicts the trend of paper counts for each year in our search corpus from 2004 to 2024.
Refer to caption Bar chart showing the number of search results returned by three databases, ACM Digital Library, IEEE Xplore, and Web of Science, for each publication year from 2004 to 2024. The bars indicate the annual volume of papers included in the initial search corpus.
@@ 3.3. Article Selection
Refer to caption PRISMA flow diagram depicting the article selection process. The figure shows the number of records identified from ACM Digital Library, IEEE Xplore, and Web of Science, followed by stages of duplicate removal, title and abstract screening, full-text eligibility assessment, and the final inclusion of studies in the review.
@@ 3.4. Data Extraction and Synthesis
Refer to caption Bar chart showing the yearly distribution of the 62 included articles across three databases, ACM Digital Library, IEEE Xplore, and Web of Science, from 2004 to 2024. Bars represent publication years, with different colors indicating contributions from each database.
@@ 3.5. Thematic Analysis
We conducted a thematic analysis to identify cohesive themes that captured patterns across the final corpus (Braun and Clarke, 2006). The goal of this analysis was to inductively derive themes grounded in the data while minimizing confirmation bias. The coder re-reviewed the articles, including their extracted data and the notes summarizing the main contributions, to generate initial codes. These codes were developed inductively, without reference to pre-existing coding frameworks or theoretical models (Nowell et al., 2017; Braun and Clarke, 2006). The codes were not mutually exclusive since the articles could address multiple codes. The coder iterated through the initial codes multiple times, conducting rounds of analysis until no new ideas emerged. Following this, the refined codes were grouped into themes, which were further revised and renamed over several rounds to improve their clarity and alignment with the corpus data. To enhance reliability and interpretive consensus, all the authors reviewed and discussed the themes and their conceptual boundaries in multiple meetings. Ultimately, the final themes were organized into three overarching categories and nine subcategories that captured the breadth and nuance of how Social VR has been studied in the context of collaboration.
@@ 4. Results
After applying the methodology described above, 62 of the 2,035 papers were included in our final corpus. Fig.  3 shows the filtering process of the papers in the IEEE Xplore Library, ACM Digital Library, and Web of Science databases. The most common reason for excluding articles during the eligibility stage was that they were not related to collaboration in VR—meaning the study either focused on collaboration without using VR (e.g., in 2D platforms), used VR without addressing collaboration, or addressed neither. This exclusion category is labeled in Fig.  3 as ”not related to VR, collaboration, or both” (n = 356). Following this process, Fig. 4 depicts the trend in the number of papers for each year among the 62 included articles from 2004 to 2024.
@@ 4.1. Description of the Included Articles
Table 1 provides an overview of research and experimental characteristics of the 62 included articles. Table 1 summarizes the research types, evaluation methodologies, and analysis methods. The majority were laboratory experiments (69.35%), followed by non-empirical papers (12.90%), survey papers (8.07%), and qualitative descriptive studies (4.84%). Three articles (4.84%) fell under other categories, including one case study, one conceptual framework, and one system architecture description.
Regarding evaluation methodology, 33.87% of the studies employed mixed methods, 32.26% used quantitative methods, and 17.74% used qualitative approaches. In 16.13% of the articles, an evaluation methodology was not employed or specified. Among the qualitative methods articles, thematic analysis was the most common method (19.36%), followed by free-text responses (4.84%) and participant experience summaries (3.23%). A subset of studies employed methods such as open coding, observational techniques, or participant experience summaries. On the quantitative side, ANOVA was the most frequently used analysis (19.28%), followed by descriptive statistics (15.67%), t-tests and Wilcoxon tests (12.05% each), and post-hoc or Mann-Whitney U tests (combined 13.24%). Many studies used more than one quantitative method, so the total count of methods across all papers exceeds 43.
@@
In terms of types of tasks, problem-solving was the most common task employed in these articles (23.26%), followed by navigation/ wayfinding (16.28%) and design/construction (16.28%). Other tasks included negotiation/planning (11.62%), meetings/discussions and data analysis (9.30% each), scientific tasks (6.98%), and one workplace simulation (2.33%). Two studies (4.65%) included task types that did not fall under the primary categories outlined in the table.
Lastly, most articles with laboratory studies examined very small teams (two or three people), which means the findings primarily reflect how small groups work together rather than how larger, more mixed teams might collaborate. Most lab studies also used convenience samples, typically university students. A smaller set recruited professionals (e.g., scientists, designers) for domain-specific tasks, and very few observed real workplace teams doing their everyday work. Reporting on recruitment locale and demographics was often limited.
@@
Table summarizing characteristics of the 62 included studies. The table lists research types, evaluation methodologies, and qualitative and quantitative analysis methods across all studies, as well as VR device types, experimental task types, and participant group sizes for the subset of laboratory studies.
@@ 4.2. Thematic Analysis
The first theme includes articles that address the foundational aspects of individual perception and experience within Social VR. This theme includes two sub-themes: “Am I Really There?,” which highlights how Social VR can simulate users’ physical presence, enhancing their interaction and engagement with the group task. The second sub-theme, “How Am I Perceived Here?,” explores avatar design and its impact on collaborative dynamics and engagement, emphasizing the importance of avatar customization to shape user interactions and self-expression within virtual spaces.
@@
The final theme examines how collaboration is supported by the unique affordances of Social VR. These affordances refer to the technical elements and interactions designed to enable realistic and functional 3D virtual environments. This theme includes three sub-themes: “Creating Realistic and Functional Virtual Environments,” which addresses how the design of virtual spaces that mimic real-world settings can improve user experience; “Facilitating Interaction through Virtual Tools and Objects,” which emphasizes the importance of object manipulation and tool use in collaborative tasks; and “Supporting Presence through Spatial Audio and Movement,” which underscores how spatial audio and coordinated movement contribute to a stronger sense of presence and effective collaborative experience. A summary of these themes, sub-themes, and their respective articles can be found in Table 2.
@@
“Am I Really There?” Sense of Presence as a Foundation for Collaboration in Social VR. 4 (Olaosebikan et al., 2022), (Khojasteh and Won, 2021), (Mei et al., 2021), (Miller et al., 2021)
“How Am I Perceived Here?” Avatar Design Shaping Collaborative Dynamics and Engagement. 11 (Osborne et al., 2023), (Torro and Pirkkalainen, 2023), (Yang et al., 2024), (Yassien et al., 2021), (Kimmel et al., 2024), (Kocur et al., 2020), (Yassien et al., 2020), (Kallioniemi et al., 2017), (Shoa et al., 2023), (Li et al., 2024), (Young et al., 2015)
@@
“Are We Really Together?” Social Presence Enhancing Team Collaboration. 12 (Olaosebikan et al., 2022), (Dey et al., 2024), (Darbar et al., 2024), (Bovo et al., 2022), (Ide et al., 2020), (Adkins et al., 2024), (Mei et al., 2021), (Kimmel et al., 2023), (Wei et al., 2022), (Weissker et al., 2023), (Moharana et al., 2022),(Chastine et al., 2005)
@@
“Are We Fully Engaged?” The Importance of Social Engagement in VR Collaboration. 7 (Xu et al., 2023), (Le Tarnec et al., 2024), (Moharana et al., 2023), (Moore et al., 2019), (Gumilar et al., 2022), (Duchowski et al., 2004), (Yang et al., 2022b)
@@
Creating Realistic and Functional Virtual Environments. 9 (Mei et al., 2021), (Osborne et al., 2023), (Men and Bryan-Kinns, 2019), (Freiwald et al., 2020), (Weissker and Froehlich, 2021), (Langa et al., 2022), (Freeman et al., 2022), (Saffo et al., 2021), (Zhang et al., 2023)
Facilitating Interaction through Virtual Tools and Objects. 11 (García et al., 2008), (García et al., 2019), (Dey et al., 2024), (Osborne et al., 2023), (Roberts et al., 2005), (Roberts et al., 2004), (dos Santos et al., 2011), (Otto et al., 2006), (Nguyen et al., 2017), (Argelaguet et al., 2011), (Rasch et al., 2024)
Supporting Presence through Spatial Audio and Movement. 10 (Weissker et al., 2019), (Immohr et al., 2024b), (Immohr et al., 2024a), (Ríos et al., 2018), (Immohr et al., 2023), (Weissker, 2020), (Naef et al., 2004), (Deacon and Barthet, 2023), (Rasch et al., 2023), (Sarmiento et al., 2014)
Table listing themes and sub-themes identified through thematic analysis of Social Virtual Reality collaboration research. Each row shows a sub-theme, the number of associated articles, and corresponding references, grouped under three main theme categories: individual perception and experience, team dynamics and collaboration, and collaboration enabled through Social VR affordances.
Table 2. Themes and sub-themes identified through thematic analysis of Social VR research on collaboration. Each row presents a sub-theme, the number of articles (N) associated with it, and its references. A total of 72 entries are shown, as some papers contribute to multiple themes.